OpenStudy (cutiekawdfi):

helP find the equation of the parabola that passes through the points. (0,-7), (7,-14),(-3,-19)

OpenStudy (lochana):

hint: standard form of parabola is \[Ax^2 + Bx + c = 0\]

OpenStudy (lochana):

we have to find A,B and C.

OpenStudy (cutiekawdfi):

how do i find it

OpenStudy (lochana):

you need to plug in points and get values and this is the equation \[y = Ax^2 + Bx + C\] first point is (0,-7) \[-7 = A0 + B0 + C\]

OpenStudy (lochana):

can you find C now?

OpenStudy (lochana):

And I forgot to put "y" in my first reply. I put 0 instead. so the correct equation is \[y = Ax^2 + Bx +C\]